OpenStudy (anonymous):

Which value for x makes the sentence true? 1/11x=1/5 a 6/55 b 11/5 c 5/11 d 1/ 55

OpenStudy (anonymous):

@katlin95

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@help123please.

OpenStudy (anonymous):

@Answerthisnow

OpenStudy (anonymous):

plz help

OpenStudy (johnweldon1993):

well.....what times 1/11...would simplify into 1/5? 11 times 1 is 11 5 times 11 is 55 11/55 how many times does 11 go into 11? 1 how many times does 11 go into 55? 5 you have 1/5 :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

@johnweldon1993 that is not it plz help righ awnser

OpenStudy (johnweldon1993):

11/5 isnt correct?

OpenStudy (anonymous):

thx

OpenStudy (johnweldon1993):

or do this....you have 1/11x = 1/5 you want x by itself so multiply both sides by 11 1x (or just x) = 11/5

OpenStudy (johnweldon1993):

just another way to look at it :)